Nicole is half as old as Donald. The sum of their ages is 72.

How old is Nicole, in years?

Answer by passionn(5)

Let Nicole's age be x
Let Donald's age be 2x
x+2x=72
3x=72
divide both sides by 3
3x/3=72/3
x=24
Nicole's age is 24
Donald's age is 72-24=48